L -1 {k / s2+k2} =sinkt

1
Expert's answer
2021-09-01T11:50:30-0400
"L(\sin(kt))=\displaystyle\int_{0}^{\infin}e^{-ts}\sin(kt)dt"

"\int e^{-ts}\sin(kt)dt=-\dfrac{1}{s}e^{-ts}\sin(kt)"

"+\dfrac{k}{s}\int e^{-ts}\cos(kt)dt"

"=-\dfrac{1}{s}e^{-ts}\sin(kt)-\dfrac{k}{s^2}e^{-ts}\cos(kt)"

"-\dfrac{k^2}{s^2}\int e^{-ts}\sin(kt)dt"

Then


"(s^2+k^2)\int e^{-ts}\sin(kt)dt=-e^{-ts}(s\sin(kt)+k\cos(kt)+C_1"

"\int e^{-ts}\sin(kt)dt=-\dfrac{e^{-ts}(s\sin(kt)+k\cos(kt)}{s^2+k^2}+C"

"L(\sin(kt))=\displaystyle\int_{0}^{\infin}e^{-ts}\sin(kt)dt"

"=\lim\limits_{A\to \infin}[-\dfrac{e^{-ts}(s\sin(kt)+k\cos(kt)}{s^2+k^2}]\begin{matrix}
 A \\
 0
\end{matrix}"

"=0+\dfrac{k}{s^2+k^2}"

"L(\sin(kt))=\dfrac{k}{s^2+k^2}"

"L^{-1}(\dfrac{k}{s^2+k^2})=\sin(kt)"